OpenStudy (anonymous):

2|x+3|=18 write in equation form please, im confused.

hero (hero):

|x+3| = 18/2 -(x+3) = 9 (x+3) = 9

jimthompson5910 (jim_thompson5910):

\[\large 2|x+3|=18\] \[\large |x+3|=9\] \[\large x+3=\pm9\] \[\large x=-3\pm9\] From there, you break it up and simplify to get x=6 or x=-12

hero (hero):

Definitely using a program. No way you can type it that fast.

jimthompson5910 (jim_thompson5910):

Or if you prefer, you can type 6,-12

jimthompson5910 (jim_thompson5910):

lol I'm flattered, but that was all by hand

hero (hero):

You must have started way before me then.

jimthompson5910 (jim_thompson5910):

I type it elsewhere, then paste

hero (hero):

I knew it.

jimthompson5910 (jim_thompson5910):

I hate this editor with a passion lol

hero (hero):

What PROGRAM are you using?

jimthompson5910 (jim_thompson5910):

I just typed that with notepad++, but if I want to render it, I use winedt

OpenStudy (anonymous):

yea what are you using. haha?

hero (hero):

Oh that. Good for you.